?(?, ?) = ?? + ?
and the budget constraint ??? + ??? ≤ ?. Assume throughout that all prices and quantities arepositive and infinitely divisible.

Derive the consumer’s indirect utility function ?(∙). Then, derive the consumer’s expenditure function, e(∙), directly from ?(∙). Finally, derive the consumer’s Hicksian/compensated demand functions (denoted ? and ? , respectively) from e(∙).
Answer #1

no- du 9+1 2 С Given that :- О: 19+х Let budget constraint be px*+py Y=M du MU, = da MUY dy MRS = = MUx Px моу Py C+1) Ру Px2px v (px, py,M) - ** (**+0) MAPY (M-Py + Martha 2РЧ М+РЯ M- Pytapy 2px 2 py = SMH py) (M+ py (2 px) (2py) Indirect utilitychange in utility = V₂ V, = 18-30,25 = - 12.25 between x and y Then we First we differentiated The U with Respect to X and Y